# Embodied Instruction Following in Unknown Environments

### v1

@@ Abstract
Table 2: Comparison with different methods on the ALFRED benchmark with the AI2THOR simulator. We conduct the experiments with different numbers of samples for model adaptation.
Method Shot Num. Test Validation
Seen Unseen Seen Unseen
SR GC SR GC SR GC SR GC
HLSM 100 0.82 6.88 0.61 3.72 0.13 2.82 0.00 1.86
FILM 100 0.20 6.71 0.00 4.23 0.00 13.19 0.00 9.65
LLM-P 100 13.05 20.58 11.58 18.47 11.82 23.54 11.10 22.44
LLM-P 100 15.33 24.57 13.41 22.89 13.53 28.28 12.92 25.35
Ours 0 16.11 28.51 14.32 27.83 13.17 27.54 13.40 31.39
5 Experiments
5.1 Implementation Details
@@
Training configurations: We employed the LLaVA-7B architecture with the Vincua-1.3-7B pre-training weights for the high-level planner and the low-level controller, which is finetuned with our generated data by the LoRA strategy. For the visual encoder, we sampled 32 visual embeddings from each frontier in the semantic feature maps up to 256 tokens as scene information representation. We generated 2k instructions with three subparts (1386 target-specific short, 333 target-specific long and 332 abstract instructions) for 2509 scenes in ProcTHOR, which results in 30k groundtruth plans for training the high-level planner. We implemented the plans in ProcTHOR with $A^{*}$ algorithm to collect the expert trajectory as the groundtruth for training low-level controller. Target-specific short and long instructions mean those containing objects to be interacted (e.g. Place the egg in the bowl) for task achievements, whose number of step plan is respectively lower than 15 and not. Abstract instructions do not contain the interacted objects in the instructions (e.g. Make a simple lunch for me). We also generate 201, 67 and 152 data for each subpart as the test set. We utilized 8 NVIDIA 3090 GPUs to finetune the high-level planner and the low-level controller for an hour in the training stage. More details are provided in Appendix A.
Figure 3: Example visualization of dynamic region attention weights.
@@
Metrics: Following the ALFRED benchmark [28], we use success rates (SR), goal condition success (GC), path length and their path-length-weighted (PLW) counterparts for evaluation. SR means the ratio of the cases where the agent completely achieve the human instructions, and GC measures the ratio of objects in the state of goal achievements. PLWSR and PLWGC calculate SR and GC weighted by the expert trajectory planning step number divided by the actual execution step number, which measures the trade-off between performance and efficiency.
@@
Simulated environments: We perform extensive experiments in the ProcTHOR and AI2THOR simulators, where the step size of translation and rotation for the agent is $0.25$m and $90^{\circ}$ respectively. ProcTHOR contains 10k house-level scenes with objects from 93 categories, where the agent receives $600\times 600$ RGB-D images in the egocentric view. We divide the scenes into small-scale ([0, 10]) and large-scale ([10, 16]) ones based on the side length of the room. AI2THOR has 120 room-level scenes with 105 categories of objects, and we follow the ALFRED setting for a fair comparison with other methods [19, 31, 4]. The agent receives $300\times 300$ RGB egocentric views and employs models provided by FILM for depth estimation.
5.2 Comparison with Baselines
@@
Table 1 demonstrates the results on ProcTHOR for LLM-Planner [31], FILM [19] and our method, where our approach significantly outperforms the state-of-the-art-method LLM-Planner. Although LLM-Planner utilizes the rich commonsense embedded in LLMs to generate plans for the agent, it fails to align the pre-trained LLM with the scene information. The generated plans are usually infeasible due to the non-existence of the objects for interaction, and the re-planning module suffers from low success rate and low efficiency. On the contrary, our method construct the semantic feature maps which grounds the pre-trained multimodal LLMs to the realistic physical scene, and the unknown environment can be efficiently explored by understanding the visual clues for executable plan generation. Our approach remains leading in performance in more challenging abstract task settings and loses less than 1% of performance on large-scale scenes compared to small ones. The semantic feature map allows our approach to efficiently find the suitable target to satisfy the instruction during active exploration. Meanwhile, the leading PLWSR and PLWGC metrics verify that our low-level controller can find the target object at a lower navigation cost. Moreover, the success rate of conventional methods in the large-scale scenes is near zero, while our approach can achieve 9.59% success rate. Since the service robot is usually deployed in house-level scenes, our method is proven to be more practical.
Figure 4: All failure cases on ProcTHOR simulator.
Table 3: Effectiveness of our generated plans and exploration actions.
Method GT Normal Specific
Plan. Exp. SR PLWSR GC PLWGC Path(m)
@@
Table 4: Ablation study of different scene feature maps.
Method Normal Specific
SR PLWSR GC PLWGC Path(m)
No Map 41.29 35.03 54.25 46.54 27.59
@@
In order to evaluate the generalization ability of our method, we compare our method with the state-of-the-art EIF approaches on the ALFRED benchmark with few or no samples for finetuning as illustrated in Table 2, where our method is evaluated in zero-shot settings without finetuning the pre-trained agent. Despite leveraging the commonsense in the pre-trained multimodal LLMs, our method provides rich visual information of the scene to ground the multimodal LLM for feasible plan generation. As a result, we achieve higher success rate with shorter path length than LLM-Planner, even though we do not leverage any samples on ALFRED for finetuning. The performance improvement is less than that on ProcTHOR because ALFRED only contains room-level scenes, and simply looking around can collect information of most objects without the requirement of semantic map construction and scene exploration.
We demonstrate the quantitative results in Figure 5, where we show the step-wise plan, the exploration process and the robot implementation during a whole sequence for EIF. In the beginning, the agent is initialized in the bedroom area and selects the navigation borders outside the room for exploration, as the instruction ’making breakfast’ is irrelevant to bedrooms. During the navigation, the agent gradually knows to explore the kitchen area by observing the dining table and the counter, and it is even aware that opening the fridge may find food for breakfast due to the rich commonsense in our finetuned low-level controller. As a result, abstract instruction is achieved by serving diverse food for breakfast, where only related regions are navigated with high exploration efficiency in the unknown environment. Figure 4 illustrates the statistics of failure cases caused by different reasons. The failure mostly comes from unsuccessful navigation because of the large house-level scene, and the top reasons including ’too close to targets’ and ’fail to see closed space’ indicate that navigation algorithms should be designed with high compatibility of the subsequent manipulation.
5.3 Ablation Studies
@@
Effectiveness of the high-level planner and the low-level controller: We evaluated the variants of our method where the planner and the controller are respectively replaced with the groundtruth step-wise plans and groundtruth action sequences. Table LABEL:table:ab_gt demonstrates the results where the performance of our methods is close to that of the groundtruth, which indicates the effectiveness of our LLaVA-based planner and controller. Moreover, the performance of planners mainly influences the success rate because it is important for the agent to interact with the right objects in the correct orders, while the controller significantly impacts the path length since directly exploring the related regions enables the agent to accomplish the instruction faster.
@@
Effectiveness of the online semantic feature map: The semantic feature map provides visual information of explored regions for the planner and the controller to generate feasible plans and efficient actions, and we report the performance of different semantic maps to validate the effectiveness of our method. Table LABEL:table:ab_map demonstrates the results for the settings of no semantic maps, semantic maps with only category information, semantic feature maps without dynamic attention and our semantic feature maps. The results demonstrate that the implicit rich semantic features are necessary for effective exploration of unknown environments, and the dynamic attention also enhances the performance of the semantic feature map as it removes the information redundancy for the large house-level scenes. We also visualize the dynamic region attention when the agent builds the semantic feature map in the unknown environment as illustrated in Figure 3. For the instruction Slice the tomato for salad, the features of the kitchen area especially the tomato and the sink are considered with high attention(The green color represents greater weight), which indicates that the dynamic region attention learns relevant visual clues for feasible action generation.
6 Conclusion
@@
In this paper, we have proposed an EIF approach for unknown environments, where the agent is required to explore the environment efficiently to generate feasible action plans with existing objects to achieve human instructions. We first build a hierarchical EIF framework including a high-level planner and a low-level controller, and then build a semantic feature map with dynamic region attention to provide visual information for the planner and the controller. Extensive experiments demonstrate the effectiveness and efficiency of our framework in the house-level unknown environment.
Limitations: This work lacks real manipulation implementation and the designed navigation policy ignores the compatibility with manipulation. We will design mobile manipulation strategies for general tasks and implement the closed-loop system on real robots in the future.
Appendix A Train Details
High-level planning and low-level controller: In the supervised instruction fine-tuning stage, we reduce memory usage via DeepSpeed ZeRO-2. The learning rate for the feature mapping layer and the LLM backbone network is set to $52\times 10^{-}$ 5, and the batch size is set to 8. Fine-tuning is performed for only one epoch. Since the semantic feature maps have been constructed through CLIP, the scene visual tokens are directly fed into the mapping layer without the visual coder during the training stage.
Visual perception: We selected 100k images from the captured expert trajectories as the training set for instance segmentation model Detic[38], fine-tuning the pre-trained model with the learning rate of $1\times 10^{-4}$ and performing 180k iterations. The batch size is set to 16 and the Adam optimiser is applied.
Appendix B Inference Details
Overview: At the $i_{th}$ time, the agent surrounds to perceive the scene information and constructs semantic feature map $S_{i}$ and frontiers mask. Then, the agent will generate the $i_{th}$ high-level planning $p_{i}$ based on the scene information, user instruction $I$ and finished step-by-step planning $\{p_{k}\}_{k=1}^{i-1}$. The low-level controller generates specific interaction actions $a_{j}^{i}$, target objects $o_{j}^{i}$ and positions $l_{j}^{i}$ based on $p_{i}$, semantic feature $\{s_{m}^{i}\}$ and textual feature $f_{m}^{i}$: 1) If the $o_{j}^{i}$ is observed by the agent, $l_{j}^{i}$ will be the location recorded on the map; 2) If not be observed, $l_{j}^{i}$ will be the frontier position. During the inference process, each instruction $I$ performs up to 30 steps of high-level planning.
input : Human instruction $I$, high level planner $HP$, low-level controller $LP$, scene observation $\mathcal{O}$, maximum number of performing step $T$.
initialization: Random load into the unknown scene;
for $i\leftarrow 0$ to $T$ do
       Constructing semantic feature map $S_{i}$ via (3);
       Generate step planning $p_{i}$ based on $S_{i}$ via (1);
       if end in $p_{i}$ then
            Break
       end if
      Compute attention $w_{i}$ and update $S_{i}$ by (4);
       Generate action $\{a_{j}^{i},l_{j}^{i},o_{j}^{i}\}_{j}^{\tau_{i}}$ via (2);
       for $j\leftarrow 0$ to $\tau_{i}$ do
             Execute $a_{j}^{i}$;
       end for
end for
Algorithm 1 Inference Process
Frontier representation: We follow [35] to generate frontier masks that distinguish between known and unknown regions based on the occupancy map. Through connected component analysis, we obtain the mask of each frontier instance. We further remove frontiers with areas smaller than the threshold (150 pixels) to reduce redundancy exploration. We sample 32 visual embeddings as frontier tokens according to the frontier instance mask on the corresponding region of the feature map, while utilizing the coordinates of their centroids for the frontier text description. The specific representation is illustrated in Figure 6 (a).
Region Attention: Since the importance of the observed image for the completed instruction is different for each frame, it is desirable to assign higher weights to task-relevant visual embeddings to enable $LC$ to generate more efficient navigation exploration planning. $HP$ generates target objects that might be required to interact to complete instruction $I$ while generating $p_{i}$ and converts them into a sentence describing $L_{dec}$. However, measuring the relevance of an image to the instruction with only a single description is not discriminative enough to highlight task-relevant regions on the feature map. To this end, we add additional variants of descriptions to calibrate the relevance of images to their corresponding descriptions for exploiting the prior knowledge of the pre-trained models extensively. Specifically, we further expand $L_{dec}$ into $\overline{L_{dec}}$ and $L_{none}$ to match the input requirements of image and text alignment models such as CLIP. $\overline{L_{dec}}$ and $L_{none}$ describe the image as not containing the target objects and not containing the objects, respectively. We adopt LongCLIP [36] to retrieve the similarity between $\smash{\left\{L_{dec},\overline{L_{dec}},L_{none}\right\}}$ and RGB images as illustrated in Figure 6 (b), and consider the score of $L_{dec}$ as the attention weight.
Appendix C More Result
Influence w.r.t. navigation frontier construction: Navigation frontier means the border between the known and unknown regions, which are represented by multiple segments.
We only select the frontiers that are longer than a threshold as the candidates for agent navigation, because extremely short frontiers usually indicate corner regions that reveals uninformative information. Therefore, we can enhance the exploration efficiency significantly. Table D illustrates the success rate and path length for different thresholds. The results demonstrate that low thresholds result is redundant navigation with high path length, while high thresholds degrade the success rate because of important scene information. We set the frontier threshold to 150 pixels to achieve higher performance and navigation cost trade-off.
Qualitative results: We demonstrate more unknown environment EIF execution sequences to reflect the superiority of our approach.
Figure 7: Our approach active search for lettuce in the fridge to complete the instruction.
Figure 8: Our approach consistently completes complex instructions for preparing breakfast in large-scale unknown scenes.
Appendix D Data
Table 5: Results regarding frontier thresholds.
      Thresholds       SR       Path(m)
      70       45.77       36.50
      100       46.27       30.41
      150       45.77       23.29
      200       43.28       19.32
Training Data: Existing EIF datasets are still limited in instruction diversity and scene scale. We design a dataset synthesis framework to minimize the generation cost and increase the scale of EIF datasets, enabling agents to adapt to large-scale unknown scenes and complex tasks. Therefore, the dataset synthesis framework consists of two main stages. The first stage is to employ GPT-4 to generate extensive high-level planning with corresponding low-level actions based on prompt and scene information, then filter logical error samples with TextWorld [29]. The second stage is to execute the interactions specifically with the oracle in the simulator, grounding the generated plans and actions into the physical scene and collecting expert trajectories.
TextWorld data generation: We collect object lists contained in each scene as scene information in ProcTHOR, consisting of the location and size of each object. GPT-4 will generate task plans based on the object information and prompts. Specifically, we annotated 22 seed tasks manually to inspire GPT-4 to generate confirmed responses. Each response contains instructions, step-by-step high-level actions, and corresponding low-level actions. We further employ self-instruction [33] to ensure the diversity of instructions(The similarity filtering threshold is set to 0.9). Meanwhile, GPT-4 will generate PDDL parameters that satisfy the ALFRED benchmarks to verify the feasibility of the planning. The generated candidate samples are sent to TextWorld and check whether the task can be executed based on the PDDL parameters to ensure the quality of the training dataset.
Grounding the generated plans: The synthetic dataset that passes the PDDL check is fed into the ProcTHOR simulator for specific interactions. We collect navigation trajectories in ProcTHOR based on the planning generated in the first stage under oracle settings, which contain RGB images, depth maps, segmentation masks, and robot poses. According to the semantic feature map building approach presented in Section 4.3, we obtain real-time semantic feature maps $S_{i}$, frontier text features $\{f_{m}^{i}\}_{m}$, and semantic embedding $\{s_{m}^{i}\}_{m}$ sequences as the agents perform interactions. Based on the step-by-step planning generated by GPT-4, we split the above sequences into step-by-step instruction-following samples. As for $HP$, we feed instruction $I$, scene information $S_{i}$ and completed steps $\{p_{k}\}_{k}^{i-1}$ as prompts, expecting to generate the next step $p_{i+1}$ to be done. Meanwhile, in order to ensure the consistency of the high-level task planning, we require $HP$ to give the planning of all subsequent actions as illustrated in Figure 9 (a). For $LC$, we take the current $p_{i}$, frontier text features $\{f_{m}^{i}\}_{m}$ and semantic features $\{s_{m}^{i}\}_{m}$ as prompt, and expect $LC$ to output the action primitives executed by the agent under the oracle setting. Specifically, if the agent observes the target object, LC generates the specific target and action primitive based on the input $p_{i}$. If it does not observe, we require $LC$ to generate the closest frontier to the oracle path as the next exploration region. $LC$ training samples are as illustrated in Figure 9 (b).
Figure 9: Visualization of training samples for high-level planner and low-level controller.
Appendix E Prompt
Figure 10 briefly demonstrates the prompt words employed to inspire GPT-4 in generating the EIF dataset, which consists of the following four main parts:
System Prompt: Primarily designed to set up the GPT-4 contextual environment for generating task planning based on a virtual robot.
Action Primitive: It is applied to constrain the scope of the interaction action generated by the GPT to ensure that it is executable.
Response Format: Ensure consistency in response format to parse specific instructions, planning and actions.
PDDL Params: Record the requirements for the completion of the instruction, including the target object and its state. An interaction is only successful if the state of the target object in the scene matches the PDDL parameter record.
Figure 10: Prompt words for GPT-4 synthetic EIF dataset.

### v2

@@ Abstract
TABLE II: Ablation experimental results of exploration strategies in the task-specific short setting, where No Exp. and No Front. represent no exploration and no frontiers exploration, respectively.
Method Normal-scale Large-scale
SR PLWSR GC PLWGC Path SR PLWSR GC PLWGC Path
No Exp. 29.85 29.09 42.08 40.92 6.09 11.27 10.68 24.26 22.99 5.32
No Front. 41.29 35.03 54.25 46.54 27.59 36.76 26.91 49.35 35.51 52.38
Ours 45.77 40.75 57.88 51.14 23.29 45.09 34.41 58.21 43.13 59.11
IV-D Data Collection and Training
Data collection: The training samples for the high-level planner consist of human instruction, current completed plans, current semantic feature maps and the groundtruth plan for the next step, and those for the low-level controller include plan for the next step, textual and semantic features for current border segments and the groundtruth action sequences representing primitives, location and targets. We leverage GPT-4 and the ProcTHOR simulator to generate the large-scale dataset to train the LLaVA-based high-level planner. We annotate several seed instructions and leverage GPT-4 to generate more instructions and corresponding plans based on the object list for each scene in the ProcTHOR, where samples with logical errors are filtered with PDDL parameters [14]. We then implement the generated plans in ProcTHOR and collect the navigation trajectories, RGB-D images, object locations and robot poses as the training data. Finally, the generated samples are parsed into high-level planning samples and low-level action data.
Training: We follow the supervised fine-tuning paradigm in LLM for training the LLaVA model in high-level planner and low-level controller, where we mask out $p_{i}$ and $\{a_{j}^{i}\}_{j=1}^{\tau_{i}}$ in the $i_{th}$ step. The CE loss leveraged in the training process is represented by:
where $\boldsymbol{X}_{V}$ denotes scene feature maps and $\boldsymbol{X}_{T}$ means input text prompt tokens. $\boldsymbol{R}_{<m}$ represents the output text tokens before the $m_{th}$ token $R_{m}$ and $M$ are number of output tokens. In the training stage, we propose to construct counterfactual samples to motivate the inference ability of the foundation model on EIF. Specifically, we remove the target objects in the scene descriptions from the original samples and replace them with target objects that have similar other properties such as usage through an artificial mapping method.
V Experiments
V-A Implementation Details
@@
Training configurations: We employed the LLaVA-7B architecture with the Vincua-1.3-7B pre-training weights for the high-level planner and the low-level controller, which is finetuned with our generated data by the LoRA strategy. For the visual encoder, we sampled 32 visual embeddings from each frontier in the semantic feature maps up to 256 tokens as scene information representation. We generated 2k instructions with three subparts (1386 target-specific short, 333 target-specific long and 332 abstract instructions) for 2509 scenes in ProcTHOR, which results in 30k groundtruth plans for training the high-level planner. We implemented the plans in ProcTHOR with $A^{*}$ algorithm to collect the expert trajectory as the groundtruth for training low-level controller. Target-specific short and long instructions mean those containing objects to be interacted (e.g. Place the egg in the bowl) for task achievements, whose number of step plan is respectively lower than 15 and not. Abstract instructions do not contain the interacted objects in the instructions (e.g. Make a simple lunch for me). We also generate 201, 67 and 152 data for each subpart as the test set. We utilized 8 NVIDIA 3090 GPUs to finetune the high-level planner and the low-level controller for an hour in the training stage.
@@
Metrics: Following the ALFRED benchmark [14], we use success rates (SR), goal condition success (GC), path length and their path-length-weighted (PLW) counterparts for evaluation. SR means the ratio of the cases where the agent completely achieve the human instructions, and GC measures the ratio of objects in the state of goal achievements. PLWSR and PLWGC calculate SR and GC weighted by the expert trajectory planning step number divided by the actual execution step number, which measures the trade-off between performance and efficiency. Path represents the distance moved by the robot, which is utilized to measure the efficiency of task planning.
@@
Simulated environments: We perform extensive experiments in the ProcTHOR simulators, where the step size of translation and rotation for the agent is $0.25$m and $90^{\circ}$ respectively. ProcTHOR contains 10k house-level scenes with objects from 93 categories, where the agent receives $600\times 600$ RGB-D images in the egocentric view. We divide the scenes into normal-scale ([0, 10]) and large-scale ([10, 16]) ones based on the side length of the room.
V-B Comparison with Baselines
@@
Table I demonstrates the results on ProcTHOR for LLM-Planner, FILM and our method, where our approach significantly outperforms the state-of-the-art-method LLM-Planner. Although LLM-Planner utilizes the rich commonsense embedded in LLMs to generate plans for the agent, it fails to align the pre-trained LLMs with the scene information. The generated plans are usually infeasible due to the non-existence of the objects for interaction, and the re-planning module suffers from low success rate and low efficiency. On the contrary, our method construct the semantic feature maps which grounds the pre-trained multimodal LLMs to the realistic physical scene, and the unknown environment can be efficiently explored by understanding the visual clues for executable plan generation. In the target-specific short task setting, it is observed that our method outperforms LLM-Planner and FILM by 17.41% and 39.80% success rate in normal scale scenes, respectively. It is worth noting that our method loses less than 2% success rate in transferring to large-scale scenes, while LLM-Planner and FILM lose 34% and 91% success rate, respectively, which demonstrates the excellent scalability of our method in scene scales. Our approach remains leading in performance in more challenging target-specific long and abstract tasks. Meanwhile, the leading PLWSR and PLWGC metrics verify that our low-level controller can find the target object at a lower navigation cost. Moreover, the success rate of conventional methods (e.g., FILM) in the large-scale scenes is near zero, while our approach can achieve 9.59% success rate. Since the service robot is usually deployed in house-level scenes, our method is proven to be more practical.
We demonstrate the qualitative results in Fig. 4, where we show the step-wise plan, the exploration process and the robot implementation during a whole sequence for EIF. In the beginning, the agent is initialized in the bedroom area and selects the navigation borders outside the room for exploration, as the instruction making breakfast is irrelevant to bedrooms. During the navigation, the agent gradually knows to explore the kitchen area by observing the dining table and the counter, and it is even aware that opening the fridge may find food for breakfast due to the rich commonsense in our finetuned low-level controller. As a result, abstract instruction is achieved by serving diverse food for breakfast, where only related regions are navigated with high exploration efficiency in the unknown environment. Fig. 5 illustrates the statistics of failure cases caused by different reasons. The failure mostly comes from unsuccessful navigation because of the large house-level scene, and the top reasons including too close to targets and fail to see closed space indicate that navigation algorithms should be designed with high compatibility of the subsequent manipulation.
TABLE III: Effectiveness of our generated plans and exploration actions.
Method GT Normal-scale & target-specific short
Plan. Exp. SR PLWSR GC PLWGC Path
@@
TABLE IV: Ablation study of different scene feature maps.
Method Normal-scale & target-specific short
SR PLWSR GC PLWGC Path
No Feature 36.36 29.53 45.73 38.06 16.46
@@
V-C Ablation Studies
@@
Effectiveness of the high-level planner and the low-level controller: We evaluated the variants of our method where the planner and the controller are respectively replaced with the groundtruth step-wise plans and groundtruth action sequences. It is important to note that some of the failure causes (e.g., too close to the target) illustrated in Fig. 5 could not be resolved even with GT step-by-step planning and navigation goals. Table LABEL:table:ab_gt demonstrates the results where the performance of our methods is close to that of the groundtruth, which indicates the effectiveness of our LLaVA-based planner and controller. Moreover, the performance of active exploration in low-level controller mainly influences the success rate, since it is important to find the correct objects to interact in unknown environments. Meanwhile, low-level controller significantly impacts the path length since directly exploring the related regions enables the agent to accomplish the instruction faster.
Figure 5: All failure cases on ProcTHOR simulator.
@@
Effectiveness of the online semantic feature map: The semantic feature map provides visual information of explored regions for the planner and the controller to generate feasible plans and efficient actions, and we report the performance of different semantic maps to validate the effectiveness of our method. Table LABEL:table:ab_map demonstrates the results for the settings of semantic maps with only category information, semantic feature maps without dynamic attention, semantic feature maps with random attention, and our semantic feature maps. The results demonstrate that the implicit rich semantic features are necessary for effective exploration of unknown environments, and the dynamic attention also enhances the performance of the semantic feature map as it removes the information redundancy for the large house-level scenes. We also visualize the dynamic region attention when the agent builds the semantic feature map in the unknown environment as illustrated in Fig. 6. For the instruction Slice the tomato for salad, the features of the kitchen area especially the tomato and the sink are considered with high attention(The green color represents greater weight), which indicates that the dynamic region attention learns relevant visual clues for feasible action generation.
TABLE V: Ablation experiment for high-level task planner.
Method Normal-scale & target-specific short
SR PLWSR GC PLWGC Path
FILM 5.97 5.97 11.17 11.17 16.55
Ours w/ BERT 24.38 18.38 39.81 29.61 20.64
Ours 45.77 40.75 57.88 51.14 23.29
TABLE VI: Ablation experiment for foundation models.
Method Normal-scale & target-specific short
SR PLWSR GC PLWGC Path
GPT-4o 30.54 19.04 43.24 27.34 20.11
Conv-LLaVA-7B 44.33 26.31 53.62 31.87 24.21
LLaMA-Adapter-7B 43.84 26.88 52.61 32.64 25.34
Ours 45.77 40.75 57.88 51.14 23.29
Effectiveness of active exploration: Existing EIF frameworks often lack active exploration capabilities, making them difficult to deploy in unknown environments. Our approach addresses this limitation by utilizing pre-trained models to construct fine-grained semantic feature maps and leveraging foundation models to generate task planning and interaction actions based on these maps. Table II demonstrates the ablation experiments for different exploration strategies in the target-specific short setting. In house-level unknown environments, the no-exploration strategy reduces success rates by 15.92% and 33.82% for normal and large-scale settings, respectively, highlighting the importance of active exploration in unknown environment EIF tasks. The efficiency of active frontier exploration is demonstrated by the fact that the success rate of the navigation strategy without frontier exploration is reduced by 4.48% and 8.30%, respectively, with comparable navigation costs compared to our approach.
Influence w.r.t. high level planner: To further clarify the performance improvement of the model, we follow the FILM setting and use BERT to recognize the target objects from the instructions and generate high-level plans by filling the target objects into the corresponding parsing templates according to the predicted task categories. Table LABEL:table:ab_exp_bert illustrates the results demonstrating that changing the LLaVA-7B to BERT occurred with performance decreases, and the performance still outperforms the FILM due to the ability of the low-level controller to explore unknown regions to find the target.
Figure 6: visualization of dynamic region attention weights.
Other foundation models: We further discuss the other foundation models and the results are illustrated in Table LABEL:table:ab_exp_mllm. GPT-4o achieved only a 30.54% success rate, significantly lower than fine-tuned foundation models. Prompt-based grounding for downstream EIF tasks is ineffective compared to fine-tuning foundation models, indicating synthesized data requires post-processing before training.
VI Conclusion
@@
In this paper, we have proposed an EIF approach for unknown environments, where the agent is required to explore the environment efficiently to generate feasible action plans with existing objects to achieve human instructions. We first build a hierarchical EIF framework including a high-level planner and a low-level controller, and then build a semantic feature map with dynamic region attention to provide visual information for the planner and the controller. Extensive experiments demonstrate the effectiveness and efficiency of our framework in the house-level unknown environment. However, this work lacks real manipulation implementation and the designed navigation policy ignores the compatibility with manipulation. We will design mobile manipulation strategies for general tasks and implement the closed-loop system on real robots in the future.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62376032, Grant 62125603 and A*STAR National Robotics Programme (RIE2025) M25N4N2009.